Entries for Issuing Bonds and Amortizing Discount by Straight-Line Method

On the first day of its fiscal year, Chin Company issued $10,600,000 of five-year, 5% bonds to finance its operations of producing and selling home improvement products. Interest is payable semiannually. The bonds were issued at a market (effective) interest rate of 6%, resulting in Chin receiving cash of $10,147,857.

a. Journalize the entries to record the following:

  1. Issuance of the bonds.
  2. First semiannual interest payment. The bond discount is combined with the semiannual interest payment. (Round your answer to the nearest dollar.)
  3. Second semiannual interest payment. The bond discount is combined with the semiannual interest payment. (Round your answer to the nearest dollar.)

If an amount box does not require an entry, leave it blank.

1. Cash
Discount on Bonds Payable
Bonds Payable
2. Interest Expense
Discount on Bonds Payable
Cash
3. Interest Expense
Discount on Bonds Payable
Cash

b. Determine the amount of the bond interest expense for the first year.
$
Answer #1
a
1 Cash 10147857
Discount on Bonds Payable   452143
      Bonds Payable 10600000
2 Interest Expense 310214
       Discount on Bonds Payable 45214 =452143/10
       Cash 265000 =10600000*5%/2
3 Interest Expense 310214
       Discount on Bonds Payable 45214
       Cash 265000
b Bond interest expense for first year 620428 =310214+310214